OpenStudy (anonymous):

solve this radical equation. 8-x=square root 2x+19

OpenStudy (anonymous):

how can you get rid of the radical sign?

OpenStudy (anonymous):

square both sides

OpenStudy (anonymous):

x=-5

OpenStudy (anonymous):

√(2x + 19) - 8 = x Add both sides by 8: √(2x + 19) = x + 8 Square both sides: √(2x + 19)² = (x + 8)² 2x + 19 = x² + 16x + 64 Bring 2x + 19 to the right: 0 = x² + 16x + 64 - (2x + 19) 0 = x² + 16x + 64 - 2x - 19 0 = x² + 14x + 45 Factoring gives: 0 = (x + 9)(x + 5) Setting each factor to zero, known as zero-product identity, we finally obtain: x + 9 = 0 and x + 5 = 0 x = {-9,-5} Check each x value to determine which one best fits the given equation: √(2(-9) + 19) - 8 = -9 √(-18 + 19) - 8 = -9 √(1) - 8 = -9 1 - 8 = -9 (FALSE) √(2(-5) + 19) - 8 = -5 √(-10 + 19) - 8 = -5 √(9) - 8 = -5 3 - 8 = -5 (TRUE) Hence, x = -5 I hope this helps!

OpenStudy (anonymous):

thanks chaser71 and jacob1233

OpenStudy (anonymous):

np

OpenStudy (anonymous):

no problem!